# Emergence of non-ergodic multifractal quantum states in geometrical fractals

### v3

@@ Abstract
Eigenstate multifractality, a hallmark of non-interacting disordered metals which can potentially be observed in many-body localized states as well, is characterized by anomalous slow dynamics and appears relevant for many areas of quantum physics from measurement-driven systems to superconductivity. We propose a novel approach to achieve non-ergodic multifractal (NEM) states without disorder by iteratively introducing defects into a crystal lattice, reshaping it from a plain structure into fractal geometry. By comprehensive analysis of the Sierpiński gasket case, we find a robust evidence of the emergence of NEM states that go beyond the conventional classification of quantum states and designate new pathways for quantum transport studies. We discuss potential experimental signatures of these states.
@@ Introduction
Since the discovery of Anderson’s localization paradigm [1], there has been a significant interest in critical phenomena that transcend the traditional dichotomy of extended and localized states. Among these, a class of novel states, commonly referred to as non-ergodic multifractal states (NEMs) [2, 3], has attracted particular attention. NEMs occupy a finite volume of the system, yet their support set, in the thermodynamic limit, has zero measure, challenging the conventional ideas of how a quantum state can be distributed in space. The concept of multifractality finds broad applications in a wide range of domains of science and engineering, including fluid dynamics [4], topography [5], finance [6], physiology [7], and epidemic modeling [8]. It also inspired advancements in machine learning, granting unexplored settings to incorporate greater hierarchical depth and achieve stable and robust prediction performances [9, 10]. In condensed matter physics, the concept of NEM is strictly connected to the singular continuous nature of the spectrum [11, 12, 13] and the critical scaling behavior [2, 14] observed at phase transitions, which give rise to anomalous subdiffusive dynamics [15, 16] and make NEM particularly intriguing in the context of quantum transport. Additionally, they exhibit sublinear growth of entanglement entropy [17, 18], defying the standard formulation of the eigenstate thermalization hypothesis [19]. These attributes are essential for understanding quantum coherence of eigenstates [20] and many-body localization phenomena [21, 22]. Moreover, it is conjectured that multifractality can enhance electron-electron interactions, heading to elevated critical temperature in superconductors [23, 24].
@@
Incommensurate potentials in one-dimensional systems [25, 26] and random disorder in higher-dimensional systems [27, 28, 29] are established mechanisms for inducing NEMs. Experimental investigation of these routes, however, often faces notable hurdles. For instance, with incommensurate systems, achieving ideal mathematical conditions is difficult; practical realizations frequently employ rational approximations whose periodicity can introduce finite-size effects or mini-gaps, potentially masking true multifractal signatures [30, 26]. Systems relying on random potentials, on the other hand, demand precise critical tuning of disorder [31]. Moreover, observations from single experimental instances may not capture the true ensemble-averaged multifractal characteristics, leading to considerable sample-to-sample variability. Beyond these specific issues, a more general challenge is that the delicate quantum interference essential for NEMs is highly vulnerable: it can be readily destroyed by slight deviations from ideal parameters, interparticle interactions, or other uncontrolled perturbations. To address these limitations, we propose an alternative approach: generating multifractal states through precise geometric manipulation, by arranging substitutional point-like defects to transform a regular lattice into a deterministic fractal geometry [32]. In recent years, properties of artificial quantum fractal structures have been systematically explored both experimentally and theoretically [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. Here, our aim is to demonstrate how the self-similar structure and non-integer dimensionality of fractals offer a natural framework for designing and studying multifractal eigenstates.
@@
We consider single-particle quantum mechanics of a tight-binding model on a Sierpiński Gasket (SG), and through a comprehensive numerical analysis, we demonstrate multiple crossovers between extended states and distinct stages of multifractality in the model spectrum. By extrapolating the fractal dimensions, we identify energy intervals where transitions to multifractal behavior occur. Examination of the singularity spectrum confirms the multifractal nature of these individual states and unveils a variety of behaviors, including NEMs that may retain metallic or insulating features. Recent advancements in spin-resolved scanning tunneling microscopy (STM) and spectroscopy techniques [48, 49] have enabled precise control over quantum states in low-dimensional structures with strong quantum confinement. Having this in mind, we also present a detailed numerical analysis of the density-density correlation functions, which can be directly observed in such experiments. By pinpointing the regions where crossovers to NEMs occur, we compute these correlators and show that they exhibit the predicted power-law decay. Overall, these attributes altogether form a robust criterion to unambiguously distinguish NEMs from both extended and localized states.
@@
While this study is of a purely fundamental theoretical nature, one aim to exploit the unconventional properties of NEMs in contexts like quantum control or superconductivity. Hence, a key question here is: can multifractal states generated with geometrical fractals, demonstrate sufficient resilience to perturbations to support their practical use? To address this, we introduce additional randomly distributed geometric defects into the system. Our analysis shows that, qualitatively, NEMs exhibit robustness against such perturbations. What is even more compelling, some NEMs react to manipulation of individual atoms in a very structured and symmetric way, possibly opening an opportunity to use them for delicate control in quantum systems.
@@
Figure 1: Transformation of regular geometry into a fractal structure: from triangular lattice (left) into a sixth generation Sierpiński gasket (SG, right). Zoom-in: the elementary building block of the SG with a substitutional defect (yellow dot). The hopping between regular sites is $t$. The defect has on-site potential $V$, and the corresponding hoppings connecting it with the rest of the lattice are $\bar{t}\ll t\ll V$.
@@
Our starting point is a uniform triangular lattice $\mathcal{L}$ containing $N_{\mathcal{L}}$ sites, each uniquely indexed. The corresponding tiaght-binding Hamiltonian is an $N_{\mathcal{L}}\times N_{\mathcal{L}}$ matrix. Due to the uniform geometry of the lattice and assuming hopping exclusively between between nearest neighbors, we set the off-diagonal elements of the Hamiltonian to $t_{ij}=t=1$ for neighboring sites $i$ and $j$ and are zero otherwise. The diagonal elements are initially set to zero, neglecting any on-site potential at this stage. The effects of including next-nearest-neighbor (NNN) hopping are discussed in the Supplemental Material (SM, sec G).
The transformation of the triangular lattice into a Sierpiński Gasket (SG) is achieved by iterative placement of defects into regions corresponding to voids in the fractal structure up to the desired generation $g$, as shown in Fig. (1). The set of defect sites forms a subset of the original lattice denoted as $\mathcal{B}\subset\mathcal{L}$, containing $N_{\mathcal{B}}$ sites. The remaining $N_{g}=N_{\mathcal{L}}-N_{\mathcal{B}}$ sites constitute the active sites of the $g$-th generation SG. Details on the site counts ($N_{\mathcal{L}}$, $N_{\mathcal{B}}$, $N_{g}$) for the generations studied are provided in the Supplemental Material (SM, see Sec. A). Each defect, located at position $r_{i}$, is modeled as a strong, localized potential: $V_{i}=V\,\delta(r-r_{i})$, with strength $V$ significantly larger than any other energy scales ($V\gg t$). To effectively isolate these defect sites and mimic the voids of the fractal, the hopping terms involving links connected to any defect site are drastically reduced ($\bar{t}\ll t$). The resulting single-particle Hamiltonian reads:
[…]
@@
For the results presented here, the model parameters are set to $\bar{t}=10^{-9}$, $t=1$, and $V=10^{4}$. We emphasize that this choice is representative, as any set of values satisfying the inequality $\bar{t}\ll t\ll V$ is sufficient to reproduce the same qualitative results for the studied properties.
@@ Results
Here, $N$ is the total number of available sites of the SG (i.e. $N=N_{g}$), and $q$ is a real parameter, referred to as the moment order. While Eq.(2) is basis-dependent, in what follows we will employ basis of localized states, with $|i\rangle$ wavefunction being fully concentrated at site $i$.
@@
The exponents $\tau_{q}$, known as mass exponents, are related to the generalized fractal dimensions $D_{q}$ by $\tau_{q}=D_{q}(q-1)$. These dimensions describe the spatial extent of the eigenstates and distinguish between different classes of quantum states. Extended (metallic-like) states spread across the entire lattice; their wave function intensity at any site $i$ typically scales as $|\langle i|\Phi\rangle|^{2}\sim 1/N$. This leads to $\text{IPR}_{q}\sim N\cdot(1/N)^{q}=N^{1-q}$, which implies $\tau_{q}=q-1$ and consequently $D_{q}=1$ for all $q\neq 1$. Conversely, localized (insulating) states occupy a finite region, $\mathcal{O}(1)$ sites, irrespective of the total system size $N$. As a result, $\text{IPR}_{q}$ approaches a constant value in the thermodynamic limit, yielding $\tau_{q}=0$ and thus $D_{q}=0$ for all $q>0$. Multifractal states represent an intermediate scenario between these two extremes. These states are non-ergodic, populating only a fractal subset of the available sites. They are characterized by mass exponents $10<\tau_{q}<q-$ 1 (or equivalently: $10<D_{q}<$ 1) and exhibit a non-trivial dependence on the moment order $q$.
@@
The moment order $q=2$ in Eq. (2) is a widely used quantity to assess the localization of eigenstates. To facilitate comparisons and visualization, we introduce a normalized version [51], $\text{nIPR}_{2}(\Phi)$, that reads as:
$\text{nIPR}_{2}(\Phi)=\frac{1}{N\cdot\text{IPR}_{2}(\Phi)}=\frac{1}{N\sum_{i=1}^{N}|\langle i|\Phi\rangle|^{4}}$ (4)
@@
In the thermodynamic limit, $\text{nIPR}_{2}$ approaches $1$ for extended states ($\text{IPR}_{2}\sim N^{-1}$) and 00 for localized states ($\text{IPR}_{2}\sim\mathcal{O}(1)$). We apply this analysis to the eigenstates of the SG tight-binding Hamiltonian (Eq. (1)), and Fig. 2 shows the computed $\text{nIPR}_{2}$ values for all such states from the $g=6$ generation, plotted against their corresponding energies.
@@
Figure 2: Normalized Inverse Participation Ratio ($\text{nIPR}_{2}$) for eigenstates of the sixth-generation Sierpiński Gasket. States are plotted by energy $E$ versus state index. The color and marker size denote $\text{nIPR}_{2}$: high values (dark red, large circles) are indicative of extended states, and low values (dark blue, small circles) of localized states. The spectrum shows the presence of different regimes; with abundance of critical states.
@@
For the aforementioned case, the $\text{nIPR}_{2}$ distribution reveals a rich energy-dependent localization landscape. For example, the ground state and several low-energy excited states exhibit high $\text{nIPR}_{2}$ values, indicative of a possible extended nature within this finite system. Localization is suggested for certain energies – as $E\simeq-2.184$ and $E=1.079$, where eigenstates have $\text{nIPR}_{2}\simeq 0$. A considerable fraction of eigenstates possesses intermediate $\text{nIPR}_{2}$ values – spanning the $(0,1)$ interval. These hint to complex, non-uniform spatial distributions that deviate from simple extended or localized forms, a characteristic feature of critical states. Such behavior is particularly anticipated in regions of high spectral density or near band gaps, where the inherent self-similar geometry of the fractal is playing a key role in shaping the eigenstate structure and support multifractality [52].
@@
Based on the $\text{nIPR}_{2}$ overview of the $g=6$ generation (Fig. 2), we select three representative energies to carry out a detailed finite-size scaling on the corresponding states, aiming to extrapolate their behavior to the thermodynamic limit. These energies are associated with: the ground state (GS), which has a high $\text{nIPR}_{2}$ value; eigenstates at $E\simeq-2.184$, which exhibit the lowest $\text{nIPR}_{2}$ values; and eigenstates at $E\simeq 1.509$, chosen from a spectral region with intermediate $\text{nIPR}_{2}$ values as potential multifractal candidates.
Our primary goal is to determine the mass exponents, $\tau_{q}(E)$, and the generalized fractal dimensions, $D_{q}(E)$, for these selected state energies. This requires analyzing the scaling of the $\text{IPR}_{q}$ with system size $N$, as described by Eq. (3) for different values of $q$.
To achieve this, we first compute $\text{IPR}_{q}(\Phi_{k})$ values using Eq. (2) for individual eigenstates $\Phi_{k}$ corresponding to the selected energies. These are obtained from SGs of varying generations, from $g=5$ to $g=10$ — corresponding to different system sizes $N$, as in SM Sec. A.
@@
To determine the exponent $\tau_{q}(E)$ in the thermodynamic limit, these finite-size values $\tau_{q}(E,N)$ are then extrapolated by plotting them against $1/\ln N$. The thermodynamic value $\tau_{q}(E)$ – and consequently the fractal dimension $D_{q}(E)=\tau_{q}(E)/(q-1)$ – is obtained as the intercept of a linear fit to these data in the limit $1/\ln N\to 0$. The outcome of this scheme is illustrated in Fig. 3.
@@
Figure 3: Finite-size scaling extrapolation of mass exponents for representative eigenstates. The panels show the effective mass exponents $\tau_{q}(E,N)$ plotted against $1/\ln N$ for various moment orders $q$. The thermodynamic limit value, $\tau_{q}(E)$, is obtained as the $y$-intercept ($1/\ln N\to 0$) of a linear fit (dashed lines) to the data points from different system sizes $N$ corresponding to Sierpiński Gasket generations $g=5$ through $g=10$.
(a) The ground state (GS), for which the extrapolated mass exponents $\tau_{q}(\text{GS})$ lead to $D_{q}(\text{GS})=\tau_{q}(\text{GS})/(q-1)\simeq 1$, confirming the extended nature of the ground state. (b) The NEM state, considered as a candidate multifractal state. The exponents $\tau_{q}$ display a distinct non-linear dependence on $q$. (c) The candidate localized state (LOC): as the extrapolated $\tau_{q}$ are found to be approximately 00 for all $q>0$.
@@
Figure 4: Extrapolated mass exponents $\tau_{q}(E)$ as a function of moment order $q$ for representative states on the Sierpiński Gasket: Ground State (GS, $E\simeq-3.667$, red circles), Localized State (LOC, $E\simeq-2.184$, blue triangles), and Non-Ergodic Multifractal State (NEM, $E\simeq 1.509$, orange squares). Statistical errors are included; where not visible, the error bars are smaller than the marker size. The dashed black line indicates the behavior of ideal extended states ($\tau_{q}=q-1$, Gaussian Orthogonal Ensemble), while the dotted black line represents ideal localized states ($\tau_{q}=0$).
@@
For the ground state (GS, $E\simeq-3.667$, red circles), the mass exponents $\tau_{q}(\text{GS})$ exhibit a clear linear dependence on $q$. This behavior closely follows the $\tau_{q}=q-1$ relation (dashed black line), characteristic of extended states, such as those in systems with Gaussian Orthogonal Ensemble (GOE) symmetry. The corresponding generalized fractal dimension $D_{2}(\text{GS})\simeq 0.997$ supporting its metallic-like nature. In contrast, the state at $E\simeq-2.184$ (LOC, blue triangles) exhibits mass exponents $\tau_{q}(\text{LOC})$ that are approximately zero across the range of positive $q$ values. This $\tau_{q}\simeq 0$ behavior (dotted black line) is representative of ideal localization, yielding generalized fractal dimensions $D_{q}(\text{LOC})\simeq 0$ for $q>0$. The specific value $D_{2}(\text{LOC})\simeq 0.017$ quantifies this strong confinement. The state at $E\simeq 1.509$ (NEM, orange squares) deviates significantly from both extended and localized limits. Its mass exponents $\tau_{q}(\text{NEM})$ display a distinct non-trivial dependence on $q$, a signature of multifractality. The corresponding fractal dimension $D_{2}(\text{NEM})\simeq 0.652$ is consistent with this multifractal character. The collective analysis of these representative states, summarized by their distinct $\tau_{q}(E)$ signatures presented in Fig. 4, unequivocally demonstrates the coexistence of extended, localized, and multifractal eigenstates within the electronic spectrum of the SG.
@@
To further validate these findings, we examine the distribution of the singularity spectrum $f(\alpha)$ [2], related via Legendre transform to the extrapolated generalized fractal dimensions $D_{q}$ (see Methods). The left panel of Fig. 5 shows the singularity spectra for the chosen representatives. The ground state (GS, red circles) exhibits a narrow $f(\alpha)$ curve sharply peaked near $\alpha_{0}=1$. This profile is in agreement with the theoretical expectation for fully extended states – as attested by the comparison with a realization of Gaussian Orthogonal Ensemble (GOE, thick black line). The insulating state (LOC, blue circles) satisfies the condition $f(\alpha\to 0)=0$, a typical mark of localization in the singularity spectrum analysis. Finally, the representative non-ergodic multifractal state (NEM, orange circles) is distinguished by a broader, roughly parabolic $f(\alpha)$ curve. Its apex is visibly shifted to a value greater than one, and the finite width of the spectrum quantifies the strength of multifractality. The robustness of the $f(\alpha)$ spectrum for the NEM state is also proved by its finite-size scaling, presented in the right panel of Fig. 5. The excellent agreement observed between $f(\alpha)$ curves computed for different ranges of system generations confirms the multifractal signature in the thermodynamic limit.
@@
Figure 5: Singularity spectrum $f(\alpha)$ for representative states and its finite-size scaling.

### v4

@@ Abstract
Eigenstate multifractality, a hallmark of non-interacting disordered metals, which may also be observed in many-body localized states, is characterized by anomalous slow dynamics and appears relevant for many areas of quantum physics, from measurement-driven systems to superconductivity. We propose a novel approach to achieve non-ergodic multifractal states (NEMs) without disorder by iteratively introducing defects into a crystal lattice, reshaping it from a plain structure into a fractal geometry. By comprehensive analysis of the Sierpiński gasket case, we find robust evidence of the emergence of NEMs that go beyond the conventional classification of quantum states and designate new pathways for quantum transport studies. We discuss potential experimental signatures of these states.
@@ Introduction
Since the discovery of Anderson’s localization paradigm [1], there has been a significant interest in critical phenomena that transcend the traditional dichotomy of extended and localized states. Among these, a class of novel states, commonly referred to as non-ergodic multifractal states (NEMs) [2, 3], has attracted particular attention. NEMs occupy a finite volume of the system, yet their support set, in the thermodynamic limit, has zero measure, challenging the conventional ideas of how a quantum state can be distributed in space. The concept of multifractality finds broad applications across science and engineering, including fluid dynamics [4], topography [5], finance [6], physiology [7], and epidemic modeling [8]. It also inspired advancements in machine learning, granting unexplored settings to incorporate greater hierarchical depth and achieve stable prediction performances [9, 10]. In condensed matter physics, the concept of NEMs is strictly connected to the singular continuous nature of the spectrum [11, 12, 13] and the critical scaling behavior [2, 14] observed at phase transitions, which give rise to anomalous subdiffusive dynamics [15, 16] and make them particularly intriguing in the context of quantum transport. Additionally, they exhibit sublinear growth of entanglement entropy [17, 18], defying the standard formulation of the eigenstate thermalization hypothesis [19]. These attributes are essential for understanding quantum coherence of eigenstates [20] and many-body localization phenomena [21, 22]. Moreover, it is conjectured that multifractality can enhance electron-electron interactions, leading to elevated critical temperature in superconductors [23, 24].
@@
Incommensurate potentials in one-dimensional systems [25, 26] and random disorder in higher-dimensional systems [27, 28, 29] are established mechanisms for inducing NEMs. Experimental investigation of these routes, however, often faces notable hurdles. For instance, with incommensurate systems, achieving ideal mathematical conditions is difficult. Practical realizations frequently employ rational approximations whose periodicity can introduce finite-size effects or mini-gaps, potentially masking true multifractal signatures [30, 26]. Systems relying on random potentials, on the other hand, demand precise critical tuning of disorder [31]. Moreover, observations from single experimental instances may not capture the true ensemble-averaged multifractal characteristics, leading to considerable sample-to-sample variability. Beyond these specific issues, a more general challenge is that the delicate quantum interference essential for NEMs is highly vulnerable: it can be readily destroyed by slight deviations from ideal parameters, interparticle interactions, or other uncontrolled perturbations. To address these limitations, we propose an alternative approach: generating multifractal states through precise geometric manipulation, by arranging substitutional point-like defects to transform a regular lattice into a deterministic fractal geometry [32]. In recent years, properties of artificial quantum fractal structures have been systematically explored experimentally and theoretically [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. Here, our aim is to demonstrate how the self-similar structure and non-integer dimensionality of fractals offer a natural framework for designing and studying multifractal eigenstates.
@@
We consider single-particle quantum mechanics of a tight-binding model on a Sierpiński gasket (SG), and through a comprehensive numerical analysis, we demonstrate multiple crossovers between extended states and distinct stages of multifractality in the model spectrum. By extrapolating the fractal dimensions, we identify energy intervals where transitions to multifractal behavior occur. Examination of the singularity spectrum confirms the multifractal nature of these individual states and unveils a variety of behaviors, including NEMs that may retain metallic or insulating features. Recent advancements in spin-resolved scanning tunneling microscopy (STM) and spectroscopy techniques [48, 49] have enabled precise control over quantum states in low-dimensional structures with strong quantum confinement. With this in mind, we also present a detailed numerical analysis of the density-density correlation functions, which can be directly observed in such experiments. By pinpointing the regions where crossovers to NEMs happen, we compute these correlators and show that they exhibit the predicted power-law decay. Altogether, these attributes form a robust criterion to unambiguously distinguish NEMs from both extended and localized states.
@@
While this study is purely theoretical, one aims to exploit the unconventional properties of NEMs in contexts like quantum control or superconductivity. Hence, a key question is: can multifractal states generated with geometrical fractals demonstrate sufficient resilience to perturbations to support their practical use? To address this, we introduce additional randomly distributed geometric defects into the system. Our analysis shows that, qualitatively, NEMs exhibit robustness against such perturbations. What is even more compelling is that some NEMs react to manipulation of individual atoms in a very structured and symmetric way, possibly opening an opportunity to use them for delicate control in quantum systems.
@@
Figure 1: Transformation of regular geometry into a fractal structure: from triangular lattice (left) into a sixth generation Sierpiński gasket (SG, right). Zoom-in: the elementary building block of the SG with a substitutional defect (yellow dot). The hopping between regular sites is $t$. The defect has on-site potential $V$, and the corresponding hoppings to the rest of the lattice are $\bar{t}\ll t\ll V$.
@@
Our starting point is a uniform triangular lattice $\mathcal{L}$ containing $N_{\mathcal{L}}$ sites, each uniquely indexed. The corresponding tight-binding Hamiltonian is an $N_{\mathcal{L}}\times N_{\mathcal{L}}$ matrix. With uniform geometry and assuming hopping exclusively between nearest neighbors, off-diagonal Hamiltonian elements are $t_{ij}=t=1$ for neighboring sites $i$ and $j$ and zero otherwise. The diagonal elements are initially set to zero, neglecting any on-site potential at this stage. The effects of including next-nearest-neighbor (NNN) hopping are discussed in the Supplemental Material (SM, Sec. F.3).
The transformation of the triangular lattice into a SG is achieved by iterative placement of defects into regions corresponding to voids in the fractal structure up to the desired generation $g$, as shown in Fig. 1. The set of defect sites forms a subset of the original lattice denoted as $\mathcal{B}\subset\mathcal{L}$, containing $N_{\mathcal{B}}$ sites. The remaining $N_{g}=N_{\mathcal{L}}-N_{\mathcal{B}}$ sites constitute the active sites of the $g$-th generation SG. Details on the site counts ($_{\mathcal{L}}, NN_{\mathcal{B}}$, $N_{g}$) for the generations studied are provided in the SM, Sec. A.
Each defect, located at position $r_{i}$, is modeled as a strong, localized potential: $V_{i}=V\,\delta(r-r_{i})$, with strength $V$ significantly larger than any other energy scales ($V\gg t$). To effectively isolate these defect sites and mimic the voids of the fractal, the hopping terms involving links connected to any defect site are drastically reduced ($\bar{t}\ll t$). The resulting single-particle Hamiltonian reads:
$H=\sum\limits_{\begin{subarray}{c}\langle i,j\rangle\\ i,j\notin\mathcal{B}\end{subarray}}t\,c^{\dagger}_{i}c_{j}+\sum_{k\in\mathcal{B}}V\,c^{\dagger}_{k}c_{k}+\sum_{\begin{subarray}{c}\langle i,k\rangle\\ k\in\mathcal{B}\end{subarray}}\overline{t}\,c^{\dagger}_{i}c_{k}+\text{h.c.}$ (1)
@@
For the results presented here, the model parameters are set to $\bar{t}=10^{-9}$, $t=1$, and $V=10^{4}$. We emphasize that this choice is representative, as any set of values satisfying the inequality $\bar{t}\ll t\ll V$ is sufficient to reproduce the same qualitative results.
@@ Results
Here, $N$ is the total number of available sites of the SG (i.e. $N=N_{g}$), and $q$ is a real parameter, referred to as the moment order. While Eq.(2) is basis-dependent, in what follows we will employ the basis of localized states, with $|i\rangle$ wavefunction being fully concentrated at site $i$.
@@
The exponents $\tau_{q}$, known as mass exponents, are related to the generalized fractal dimensions $D_{q}$ by $\tau_{q}=D_{q}(q-1)$. These dimensions describe the spatial extent of the eigenstates and distinguish between different classes of quantum states. Within this framework, the dimensions are normalized to the geometry under study, ensuring that $D_{q}$ directly quantifies the fraction of sites occupied by the eigenstate [51, 52]. Extended (metallic-like) states spread across the entire lattice; their wave function intensity at any site $i$ typically scales as $|\langle i|\Phi\rangle|^{2}\sim 1/N$. This leads to $\text{IPR}_{q}\sim N\cdot(1/N)^{q}=N^{1-q}$, which implies $\tau_{q}=q-1$ and consequently $D_{q}=1$ for all $q$. Conversely, localized (insulating) states occupy a finite region, $\mathcal{O}(1)$ sites, independent of $N$. As a result, $\text{IPR}_{q}$ approaches a constant value in the thermodynamic limit, yielding $\tau_{q}=0$ and thus $D_{q}=0$ for all $q>0$. Multifractal states represent an intermediate scenario between these two extremes. These states are non-ergodic, populating only a fractal subset of the available sites. They are characterized by mass exponents $0<\tau_{q}<q-1$ (or equivalently: $0<D_{q}<1$) and exhibit non-trivial $q$-dependence.
@@
The moment order $q=2$ in Eq. (2) is a widely used quantity to assess the localization of eigenstates. To facilitate comparisons and visualization, we introduce a normalized version [53], $\text{nPR}_{2}(\Phi)$, that reads as:
$\text{nPR}_{2}(\Phi)=\frac{1}{N\cdot\text{IPR}_{2}(\Phi)}=\frac{1}{N\sum_{i=1}^{N}|\langle i|\Phi\rangle|^{4}}$ (4)
@@
In the thermodynamic limit, $\text{nPR}_{2}$ approaches $1$ for extended states ($\text{IPR}_{2}\sim N^{-1}$) and 0 for localized states ($\text{IPR}_{2}\sim\mathcal{O}(1)$). We apply this analysis to the eigenstates of the SG tight-binding Hamiltonian (Eq. (1)), and Fig. 2 shows the computed $\text{nPR}_{2}$ values for all such states from the $g=6$ generation, plotted against their corresponding energies.
@@
Figure 2: Normalized Participation Ratio ($\text{nPR}_{2}$) for eigenstates of the sixth-generation Sierpiński gasket. States are plotted by energy $E$ versus state index. The color and marker size denote $\text{nPR}_{2}$: high values (dark red, large circles) are indicative of extended states, and low values (dark blue, small circles) of localized states. Owing to the spectrum’s hierarchical structure (SM, Sec. B) distinct regimes emerge, with an abundance of critical states.
@@
For the aforementioned case, the $\text{nPR}_{2}$ distribution reveals a rich energy-dependent localization landscape. For example, the ground state and several low-energy excited states exhibit high $\text{nPR}_{2}$ values, indicative of a possible extended nature within this finite system. Localization is suggested for certain energies— at $E\simeq-2.184$ and $E\simeq 1.079$, where the eigenstates have $\text{nPR}_{2}\simeq 0$. A considerable fraction of eigenstates possesses intermediate $\text{nPR}_{2}$ values— spanning the $(0,1)$ interval. These hint at complex, non-uniform spatial distributions that deviate from simple extended or localized forms, a characteristic feature of critical states. Such behavior is particularly anticipated in high spectral density regions or near band gaps, where the fractal’s self-similar geometry shapes eigenstates and supports multifractality [54].
@@
Based on the $\text{nPR}_{2}$ overview of the $g=6$ generation (Fig. 2), we select three representative energies to carry out a detailed finite-size scaling of the corresponding states, extrapolating their behavior to the thermodynamic limit. These energies are associated with: the ground state (GS), which has a high $\text{nPR}_{2}$ value; the degenerate triplet at $E\simeq-2.184$, which exhibits the lowest $\text{nPR}_{2}$ values; and eigenstates at $E\simeq 1.509$, chosen from a spectral region with intermediate $\text{nPR}_{2}$ values as potential multifractal candidates. Our primary goal is to determine the mass exponents, $\tau_{q}(E)$, and the generalized fractal dimensions, $D_{q}(E)$. This requires analyzing the scaling of the $\text{IPR}_{q}$ with system size $N$, as described by Eq. (3) for different values of $q$. To achieve this, we first compute $\text{IPR}_{q}(\Phi_{k})$ values using Eq. (2) for individual eigenstates $\Phi_{k}$ corresponding to the selected energies. These are obtained from SGs of varying generations, from $g=5$ to g=10g=10— corresponding to different system sizes $N$, as in SM Sec. A.
@@
To determine the exponent $\tau_{q}(E)$ in the thermodynamic limit, these finite-size values $\tau_{q}(E,N)$ are then extrapolated by plotting them against $1/\ln N$. The thermodynamic value $\tau_{q}(E)$—and consequently the fractal dimension $D_{q}(E)=\tau_{q}(E)/(q-1)$— is obtained as the intercept of a linear fit to these data in the limit $1/\ln N\to 0$. The outcome is illustrated in Fig. 3.
@@
Figure 3: Finite-size scaling extrapolation of mass exponents for representative eigenstates. The panels show the effective mass exponents $\tau_{q}(E,N)$ plotted against $1/\ln N$ for various moment orders $q$. The thermodynamic limit value, $\tau_{q}(E)$, is obtained as the $y$-intercept ($1/\ln N\to 0$) of a linear fit (dashed lines) to the data points from different system sizes NN corresponding to Sierpiński gasket generations $g=5$ through $g=10$.
(a) The ground state (GS), for which the extrapolated mass exponents $\tau_{q}(\text{GS})$ lead to $D_{q}(\text{GS})=\tau_{q}(\text{GS})/(q-1)\simeq 1$, confirming the extended nature of the ground state. (b) The NEM state, considered as a candidate multifractal state. The exponents $\tau_{q}$ display a distinct non-linear dependence on $q$. (c) The candidate localized state (LOC): as the extrapolated $\tau_{q}\simeq 0$ for all $q>0$.
@@
Figure 4: Extrapolated mass exponents $\tau_{q}(E)$ as a function of moment order $q$ for representative states on the Sierpiński gasket: Ground State (GS, $E\simeq-3.667$, red circles), Localized State (LOC, $E\simeq-2.184$, blue triangles), and Non-Ergodic Multifractal State (NEM, $E\simeq 1.509$, orange squares). Statistical errors are included; where not visible, the error bars are smaller than the marker size. The dashed black line indicates the behavior of ideal extended states ($\tau_{q}=q-1$, as for Gaussian Orthogonal Ensemble), while the dotted black line represents ideal localized states ($\tau_{q}=0$).
@@
For the ground state (GS, $E\simeq-3.667$, red circles), the mass exponents $\tau_{q}(\text{GS})$ exhibit a clear linear dependence on $q$. This behavior closely follows the $\tau_{q}=q-1$ relation (dashed black line), characteristic of extended states, such as those in systems with Gaussian Orthogonal Ensemble (GOE) symmetry. The corresponding generalized fractal dimension $D_{2}(\text{GS})\simeq 0.997$, which supports its metallic-like nature. In contrast, the state at $E\simeq-2.184$ (LOC, blue triangles) exhibits mass exponents $\tau_{q}(\text{LOC})$ that are approximately zero across the range of positive $q$ values. This $\tau_{q}\simeq 0$ behavior (dotted black line) is representative of exponential localization (see also SM, Sec.C.2), yielding generalized fractal dimensions $D_{q}(\text{LOC})\simeq 0$ for $q>0$. The specific value $D_{2}(\text{LOC})\simeq 0.017$ quantifies this strong confinement. The state at $E\simeq 1.509$ (NEM, orange squares) deviates significantly from both extended and localized limits. Its mass exponents $\tau_{q}(\text{NEM})$ display a distinct non-trivial dependence on $q$, a signature of multifractality; its fractal dimension $D_{2}(\text{NEM})\simeq 0.652$ further supports this. The collective analysis of these representative states, summarized by their distinct $\tau_{q}(E)$ signatures presented in Fig. 4, unequivocally demonstrates the coexistence of extended, localized, and multifractal eigenstates within the spectrum of the SG.
@@
To further validate these findings, we examine the distribution of the singularity spectrum $f(\alpha)$ [2]. We compute the spectra shown via the Legendre transform of the extrapolated mass exponent $\tau_{q}$ (see Methods), while an alternative analysis using the histogram method is presented in the SM, Sec. C.3. The left panel of Fig. 5 compares the resulting hull of $f(\alpha)$ for the chosen representatives. The ground state (GS, red circles) exhibits a narrow $f(\alpha)$ curve sharply peaked near $\alpha_{0}=1$. This profile is in agreement with the theoretical expectation for fully extended states— as attested by the comparison with a realization of Gaussian Orthogonal Ensemble (GOE, thick black line). The insulating state (LOC, blue circles) satisfies the condition $f(\alpha\to 0)=0$, a typical mark of localization in the singularity spectrum analysis. Finally, the representative non-ergodic multifractal state (NEM, orange circles) is distinguished by a broader, roughly parabolic $f(\alpha)$ curve. Its apex is visibly shifted to a value greater than one, and the finite width of the spectrum quantifies the strength of multifractality.
The robustness of the $f(\alpha)$ spectrum for the NEM state is also proved by its finite-size scaling, presented in the right panel of Fig. 5. This anchored-window [15] analysis compares spectra from system size ranges sharing the largest available generation. The resulting spectra remain broad with substantial overlap, supporting the multifractal character of the state. On the other hand, the slow convergence, visible as a slight drift between the curves, is due to strong finite-size effects rather than a flow towards an ergodic state (see SM, Sec. C.3).
@@
Figure 5: Singularity spectrum $f(\alpha)$ for representative states and its finite-size scaling. All spectra are calculated using the same range of moment orders $q$ as presented in Fig. 4.